can the velocity of an object reverse direction when the objects acceleraion is constant

Grade:12

1 Answers

Arun
6 years ago
= 1\ m / s^{2}= \frac{d(t-2)}{dt}a (t) = \frac{dt(t)}{dt}Later as the value of increases and becomes greater than 2 sec , the sign of the velocity changes and becomes positive, that is the object start to move in the backward direction.At t= 3 sec, the velocity is:v (t) = 1While the acceleration remains:The acceleration of the object is the derivative of v (t) and is:= 1\ m/\ s^{2}= \frac{d(t-2)}{dt}a (t) = \frac{dv(t)}{dt}At t = 2 the velocity v (t) of the object is:v (t) = t – 2= 2 – 2= 0 m / sHowever, the acceleration remains constant:The acceleration of the object is the derivative of v (t) and is:= 1\ m/\ s^{2}= \frac{d(t-2)}{dt}a (t) = \frac{dv(t)}{dt}Yes, the velocity of the object can reverse direction when its acceleration is constant.For example consider that the velocity of any object at any time t is given as:v(t) = t – 2.At At t = 0 sec, the magnitude of velocity is 2m/s and is moving in the forward direction i.e.v (t) = -2. The negative sign here direction of motion, we consider the forward direction to be negative and backward to be positive.The acceleration of the object is the derivative of v(t) and is;